A copper wire has a resistance of 0.39Ω at 20.0°C, and an iron wire has a resistance of 0.6Ω at the same temperature. At what temperature are their resistances equal?


Round your answer to 2 decimal places


1
Expert's answer
2021-05-16T17:52:18-0400
"R_c(1+\alpha_c(T-T_0))=R_i(1+\alpha_i(T-T_0))\\
0.39(1+0.004041(T-20))\\=0.6(1+0.005671(T-20))\\T=-95.0\degree C"